OpenStudy (anonymous):

determing whether the ordered pair is a solution of the inequality (-5,6); 5x + y < - 3

OpenStudy (anonymous):

All we have to do is substitute (-5,6) to the inequality 5x+y<-3 x=-5, y=6 5(-5)+6<-3 -25+6<-3 -19<-3 True, so (-5,6) is a solution to the inequality.

OpenStudy (anonymous):

thanks pat18 you're awesome

OpenStudy (anonymous):

you are welcome. it is been so long since i am online.lol

OpenStudy (anonymous):

I know i missed you

OpenStudy (anonymous):

thank you. i miss all.

OpenStudy (anonymous):

pat18 can u help with this one solve the compound inequality 4x-1/4 < or equal too - 3 or 4x-1/4 >or equal too 1

OpenStudy (anonymous):

Actually, we can solve this in a step by step manner. 4x-1/4<=-3 4x<=-3+1/4 4x<=-11/4 x<=-11/8 4x-1/4>=1 4x>=1+1/4 4x>=5/4 x>=5/8 So, we can have a solution of -11/8>=x>=5/8

OpenStudy (anonymous):

the value of x is x is lesser than or equal to -11/8 but greater than or equal to 5/8

OpenStudy (anonymous):

thanks pat18 good answer